OpenStudy (anonymous):

If y=x^3+2x and dx/dt=6. Find dy/dt when x=5

OpenStudy (bahrom7893):

dy/dt = 3x^2(dx/dt)+2(dx/dt) dy/dt = 3*5^2*6 + 2*6 = 3*25*6+12 = 75*6 + 12 = 450+12 = 462

OpenStudy (anonymous):

If z^2=x^2+y^2,dx/dt=9, and dy/dt=6, find dz/dt when x=2 and y=4

OpenStudy (bahrom7893):

do the same thing Jinnie

OpenStudy (anonymous):

haha yeah

OpenStudy (bahrom7893):

2z(dz/dt)...........

OpenStudy (anonymous):

im a little confuse since there are two different d's

OpenStudy (bahrom7893):

z^2=x^2+y^2 2z*(dz/dt) = 2x(dx/dt) + 2y(dy/dt) plug in all values.

OpenStudy (anonymous):

ok understood

OpenStudy (anonymous):

can you take a look at this one, its a different concept

OpenStudy (bahrom7893):

sure

OpenStudy (anonymous):

A street light is mounted at the top of a 11 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 7 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 35 ft from the base of the pole?

OpenStudy (bahrom7893):

Typical related rates problem. And you have to use similar triangles.

OpenStudy (bahrom7893):

BE/AB = DE/CD => plug in all values that you have, and take derivative

OpenStudy (anonymous):

ok thanks are you like a math teacher?

OpenStudy (bahrom7893):

i did work as a tutor lol, but no i'm just a teenager haha

OpenStudy (anonymous):

what?!!! how old are you?

OpenStudy (anonymous):

what?!!! how old are you?

OpenStudy (bahrom7893):

18

OpenStudy (bahrom7893):

actually turning 19 in 5 months

OpenStudy (anonymous):

whoa how did you learn all this at age 18

OpenStudy (bahrom7893):

haha good question. anyway back to your homework.. That's a typical shadow woman walking away from a lamp post question lol

OpenStudy (anonymous):

im really urious im a sophmore in college and idont know all this lol

OpenStudy (bahrom7893):

im a junior in college, and i've taken calc1-3, diff equations, topology, lin alg, so i know all this lol.

OpenStudy (anonymous):

you're 18 and a junior in college?

OpenStudy (anonymous):

Wow, bahrom just 18 with all Cal stuff and topology! Look like you aim to network security major. I wish my cousin just 1/100th as mature as you :)

OpenStudy (anonymous):

y' = 3x^2 x' + 2 x' = 3 * 5^2 * 6 + 2 * 6 = 6( 75 + 2) = 6 * 77 = 462

OpenStudy (bahrom7893):

lol im a math major

OpenStudy (anonymous):

chlorophyll thanks for the help. i was focused on this question A street light is mounted at the top of a 11 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 7 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 35 ft from the base of the pole? you can scroll up and see an attachment i posted along with it

OpenStudy (anonymous):

Jinnie, it's much better for you to learn from bahrom!

OpenStudy (anonymous):

Anyway, I'm going to be back after feeding my precious kitten :)

OpenStudy (anonymous):

alright thanks for all the help

OpenStudy (anonymous):

Suppose X is the total distance ( the leg in red) in which: S is shadow of the woman, x is distance from woman to the base of street light -> Tip of the shadow: X = s + x => Rate of tip shadow moving: X' = s' + x' Given x' = +7ft/sec By ratio: s/6 = ( s + x)/ 11 (similar triangle) -> s = 6x/5 => s' = (6/5) x' = (6/5) * (7) = 42/5 Thus X' = 42/5 + 7 = (42 + 35) / 5 = 15.4 ft/sec

OpenStudy (anonymous):

I keep considering the distance 35 ft, and don't know how to apply in this situation ???